40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4032     Entry time: Thu Dec 9 00:34:53 2010
Author: Osamu 
Type: Update 
Category: SUS 
Subject: ITMX oplev Pitch OLTF 

We measured Open loop TF for oplev pitch on ITMX.

2010128_01_ITMX_oplev_pit_oltf_fine.jpg

2010128_01_ITMX_oplev_pit_oltf_fine.pdf

 All feed back filter of oplev  was on as same as before. Original notch filters which notches above 10Hz resonance should be modified with some measurements of present resonant frequency. Up to 10Hz, a simple f^2 filter is used, so the notch should not affect this measurement.

Measured upper UGF is about 2Hz with gain slider 1, and lower UGF is 1.3Hz. Phase margin is 40 degree, so it is not a good idea to increase the gain drastically.

I have measured the coherence also but I could not find a way to put it on this picture. Anyway coherence below 0.6Hz was not so good like ~0.95. This can be improved if larger excitation is used next time.

During this measurement around 0.2-0.3Hz, small earthquake happened but seemed OK for the control.

We will measure the other TF, yaw, ETMX or somthing, maybe tomorrow, due to free swinging ITMX and ETMX tonight.

 

ELOG V3.1.3-